The marginal propensity to consume in an economy is

0.75

​ Thus, if the price level is​ fixed, a

​$300 increase in investment

would be expected to increase real GDP by ​$______???

​(Round your answer to the nearest​ dollar.)
Answer #1

Answer

Multiplier =1/(1-MPC)

=1/(1-0.75)

=4

Change in GDP =change in investment * multiplier  

=300*4

=$1200

the GDP increases by $1200